# CLEF 2024 – Conference and Labs of the Evaluation Forum, September 9–12, 2024, Grenoble, France

### v1

@@ 1 Introduction
This paper describes the solution of the SmurfCat team, which achieved first place with an average score of 0.52 in the automatic evaluation and second place with a score of 0.74 in the manual human evaluation. Our solution is based on the mT0 model family [3], which has powerful multilingual capabilities. We fine-tuned all our selected models to each language of the competition, and applied various data augmentation techniques. To improve detoxification, we performed hypothesis filtering using the diverse beam search algorithm [4]. Finally, we applied ORPO [5] alignment to enforce model predictions. Our 3.7-billion-parameter language model demonstrates state-of-the-art results for Ukrainian and near state-of-the-art results for other languages. We published the final best-performing model on the HuggingFace Hub***https://hf.co/s-nlp/mt0-xl-detox-orpo. You can also find the training scripts and the extended data on GitHub***https://github.com/s-nlp/multilingual-transformer-detoxification.
@@ 2 Data
Initially, there were not many parallel datasets for the multilingual detoxification task. More precisely, primarily only the Russian***https://huggingface.co/datasets/s-nlp/ru_paradetox and English***https://huggingface.co/datasets/s-nlp/paradetox ParaDetox datasets were available, with 11 100 and 19 700 samples respectively. During the competition, the organizers published a small human-annotated Multilingual ParaDetox***https://huggingface.co/datasets/textdetox/multilingual_paradetox for all languages, containing only 400 samples per language.
Nevertheless, we decided to augment the provided data by automatic translation from English to other languages. To translate the original English data, we used a GoogleTranslator model from deep_translator***https://pypi.org/project/deep-translator/ Python package. We chose API over some of the more advanced machine translation models because of its speed and simplicity. Also, there are not as many translators for low-resource languages like Amharic. As a result, we obtained an additional 19 700 samples for each language.
@@
Since translation is often imperfect, we decided to perform a specific post-processing procedure. In general, we checked the preservation of meaning after translation and the toxicity of the translated data. First, we used the LaBSE [6] model to evaluate the similarity between translated pairs. Second, we applied XLM-R***https://huggingface.co/textdetox/xlmr-large-toxicity-classifier toxicity classifier to check whether toxic sentences were still toxic after translation and vice versa.
@@
Table 1: Statistics of number of remaining samples after filtering.

### v2

# CLEF 2024: Conference and Labs of the Evaluation Forum, September 09–12, 2024, Grenoble, France

@@ 1 Introduction
This paper describes the solution of the SmurfCat team, which achieved first place with an average score of 0.52 in the automatic evaluation and second place with a score of 0.74 in the manual human evaluation. Our solution is based on the mT0 model family [3], which has powerful multilingual capabilities. We fine-tuned all our selected models to each language of the competition, and applied various data augmentation techniques. To improve detoxification, we performed hypothesis filtering using the diverse beam search algorithm [4]. Finally, we applied ORPO [5] alignment to enforce model predictions. Our 3.7-billion-parameter language model demonstrates state-of-the-art results for Ukrainian and near state-of-the-art results for other languages. We published the final best-performing model on the HuggingFace Hub [^1] https://hf.co/s-nlp/mt0-xl-detox-orpo. You can also find the training scripts and the extended data on GitHub [^2] https://github.com/s-nlp/multilingual-transformer-detoxification.
@@ 2 Data
Initially, there were not many parallel datasets for the multilingual detoxification task. More precisely, primarily only the Russian [^3] https://huggingface.co/datasets/s-nlp/ru_paradetox and English [^4] https://huggingface.co/datasets/s-nlp/paradetox ParaDetox datasets were available, with 11 100 and 19 700 samples respectively. During the competition, the organizers published a small human-annotated Multilingual ParaDetox [^5] https://huggingface.co/datasets/textdetox/multilingual_paradetox for all languages, containing only 400 samples per language.
Nevertheless, we decided to augment the provided data by automatic translation from English to other languages. To translate the original English data, we used a GoogleTranslator model from deep_translator [^6] https://pypi.org/project/deep-translator/ Python package. We chose API over some of the more advanced machine translation models because of its speed and simplicity. Also, there are not as many translators for low-resource languages like Amharic. As a result, we obtained an additional 19 700 samples for each language.
@@
Since translation is often imperfect, we decided to perform a specific post-processing procedure. In general, we checked the preservation of meaning after translation and the toxicity of the translated data. First, we used the LaBSE [6] model to evaluate the similarity between translated pairs. Second, we applied XLM-R [^7] https://huggingface.co/textdetox/xlmr-large-toxicity-classifier toxicity classifier to check whether toxic sentences were still toxic after translation and vice versa.
@@
Table 1: Statistics of number of remaining samples after filtering.